# Krakow,Warsaw,Warsaw

### v1

@@ Abstract
The proper design and architecture of testing of machine learning models, especially in their application to quantitative finance problems, is crucial. The most important in this process is selecting an adequate loss function used for training, validation, estimation purposes, and tuning of hyperparameters. Therefore, in this research, through empirical experiments on equity and cryptocurrency assets, we introduce the Mean Absolute Directional Loss (MADL) function which is more adequate for optimizing forecast-generating models used in algorithmic investment strategies. The MADL function results are compared for Transformer and LSTM models and we show that almost in every case Transformer results are significantly better than those obtained with LSTM.
@@ 1 Introduction
The starting point of this research focuses on several key issues at the intersection of machine learning and quantitative finance. Firstly, there is a theoretical focus on determining the most suitable architecture for testing machine learning forecasting models. Secondly, it includes practical efforts to use these forecasts to generate signals for algorithmic investment strategies. Thirdly, it involves testing and comparing Transformer models with LSTM models to evaluate their effectiveness in investment strategies. Lastly, there is practical testing of empirical data from stock and cryptocurrency markets across multiple assets.
The main goal of this research is to apply the transformer model to time series forecasting, using a newly introduced loss function (MADL). We also compare the transformer with the LSTM using two types of asset classes. There are two opposing sides in the scientific community: one saying that transformers can be successfully applied to time series forecasting, and one that they can’t and shouldn’t. Both sides provide significant examples and research to prove their point. We intend to engage in this discourse and conduct our comprehensive research.
@@
The methodology is based on the application of two alternative models (Transformer and LSTM) to generate long/short signals for two types of assets: crypto (Bitcoin, Ethereum, and Litecoin) and equity (JP Morgan, S&P500 and Exxon Mobil Corp) with daily data. To keep the out-of-sample period as long as possible, a walk-forward procedure was applied. The performance of the trading strategies is evaluated using risk-adjusted returns, drawdown metrics, and equity lines.
We contribute to the literature in the following ways. First, we present the application of an adequate loss function (MADL) in ML models to generate trading signals. Second, we verify the advantages of using the transformer model over the LSTM in algorithmic trading. Third, we apply a strict methodology for six assets, controlling the overfitting effects, applying a walk-forward procedure, and extending the the out-of-sample period for 9+ years for equity and 8+ years for crypto assets.
@@ 3.5 Loss Function
If we frame the problem this way, the value of the function will be equal to the observed return on the investment with the predicted direction. This allows the model to indicate whether its prediction will result in a profit or a loss, as well as quantify the expected profit or loss. MADL was specifically designed to work with AISs rather than only verifying point forecasts. The function in our model is minimized, so that if it gives negative values, the strategy will make a profit, and if it gives positive values, the strategy will generate a loss.
@@ 5 Conclusions
In this study, we evaluate the application of the Mean Absolute Directional Loss function ([14]) in algorithmic trading with two machine learning algorithms: the transformer model ([18]) and the LSTM ([9]). The models were applied to the daily data of six assets (cryptocurrencies, including Bitcoin, Ethereum, and Litecoin, and equity stocks, including JP Morgan, S\$P 500, and Exxon Mobil). The walk-forward procedure was used to include the out-of-sample period, which was as long as 8+ years.
The results show that we successfully adapted the basic transformer model architecture to produce trading strategies yielding abnormal risk-adjusted returns. The transformer model outperforms the Buy&Hold and LSTM-based strategies for both types of assets under investigation. Transformer models produce higher risk-adjusted returns compared with both the LSTM and the Buy&Hold strategy.
Our contribution to the literature is threefold. First, we demonstrate the application of an appropriate loss function (MADL) within machine learning models to generate trading signals. Second, we assess the advantages of using transformer models over LSTM models in algorithmic trading. Third, we apply a rigorous methodology across six assets, carefully controlling for overfitting, implementing a walk-forward procedure, and extending the out-of-sample period to over nine years for equities and more than eight years for cryptocurrency assets.
The findings from this study carry several potential policy implications, particularly for financial market regulation and algorithmic trading oversight. First, the demonstrated ability of transformer models to consistently outperform traditional strategies highlights the growing role of advanced machine learning in generating high risk-adjusted returns. This might prompt regulatory bodies to consider new guidelines for algorithmic trading practices, especially regarding transparency and risk management. Furthermore, given the long out-of-sample testing period and robust methodology employed, these findings may encourage policy discussions around implementing stricter standards for the validation and monitoring of algorithmic models to safeguard against overfitting and ensure consistent performance. Finally, as these advanced models could widen the gap between retail and institutional investors, policies may be required to promote equitable access to AI-driven trading technologies.

### v2

@@ Abstract
The proper design and architecture of testing machine learning models, especially in their application to quantitative finance problems, is crucial. The most important aspect of this process is selecting an adequate loss function for training, validation, estimation purposes, and hyperparameter tuning. Therefore, in this research, through empirical experiments on equity and cryptocurrency assets, we apply the Mean Absolute Directional Loss (MADL) function, which is more adequate for optimizing forecast-generating models used in algorithmic investment strategies. The MADL function results are compared between Transformer and LSTM models, and we show that in almost every case, Transformer results are significantly better than those obtained with LSTM.
@@ 1 Introduction
This research focuses on several key issues at the intersection of machine learning and quantitative finance. First, there is a theoretical focus on determining the most suitable architecture for testing machine learning forecasting models. Second, it includes practical efforts to use these forecasts to generate signals for algorithmic investment strategies. Third, it involves testing and comparing Transformer models with LSTM models to evaluate their effectiveness in investment strategies. Finally, there is practical testing of empirical data from stock and cryptocurrency markets across multiple assets.
The main goal of this research is to apply the transformer model to time series forecasting, using a recently introduced loss function (MADL). We also compare the transformer with the LSTM using two types of asset classes. There are two opposing sides in the scientific community: one saying that transformers can be successfully applied to time series forecasting, and one that they can’t and shouldn’t. Both sides provide significant examples and research to prove their point. We intend to engage in this discourse and conduct our comprehensive research.
@@
The methodology is based on the application of two alternative models (Transformer and LSTM) to generate long/short signals for two types of assets: cryptocurrencies (Bitcoin, Ethereum, and Litecoin) and equities (JP Morgan, S&P500 and Exxon Mobil Corp) with daily data. To keep the out-of-sample period as long as possible, a walk-forward procedure was applied. The performance of the trading strategies is evaluated using risk-adjusted returns, drawdown metrics, and equity lines.
We contribute to the literature in the following ways. We present the application of an adequate loss function (MADL) in ML models to generate trading signals. Additionally, we verify the advantages of using the transformer model over the LSTM in algorithmic trading. Furthermore, we apply a strict methodology for six assets, controlling the overfitting effects, applying a walk-forward procedure, and extending the out-of-sample period for 9+ years for equity and 8+ years for cryptocurrency assets.
@@ 3.5 Loss Function
When we approach the problem from this perspective, the function’s value corresponds to the actual investment return given the forecasted direction. This allows the model to determine not only whether its prediction will yield a profit or loss, but also to estimate the magnitude of that expected gain or loss. MADL was purposefully developed for use with algorithmic investment strategies (AIS) beyond simply validating point predictions. Since our model minimizes this function, negative values indicate profitable strategy outcomes, while positive values signal potential losses.
@@ 5 Conclusions
In this study, we evaluate the application of the Mean Absolute Directional Loss function ([14]) in algorithmic trading with two machine learning algorithms: the transformer model ([18]) and the LSTM ([9]). The models were applied to daily data from six assets (cryptocurrencies, including Bitcoin, Ethereum, and Litecoin, and equity stocks, including JP Morgan, S&P 500, and Exxon Mobil). The walk-forward procedure was used to include the out-of-sample period, which extended over eight years.
The results show that we successfully adapted the basic transformer architecture to produce trading strategies yielding abnormal risk-adjusted returns. The transformer outperforms the Buy&Hold and LSTM-based strategies for both types of assets under investigation. This architecture produces higher risk-adjusted returns compared with both the LSTM and the Buy&Hold strategy.
Our contribution to the literature is threefold. First, we demonstrate the application of an appropriate loss function (MADL) within machine learning models to generate trading signals. Second, we assess the advantages of using transformer models over LSTM models in algorithmic trading. Third, we apply a rigorous methodology across six assets, carefully controlling for overfitting, implementing a walk-forward procedure, and extending the out-of-sample period to over nine years for equities and over eight years for cryptocurrency assets.
The findings from this study carry several potential policy implications, particularly for financial market regulation and algorithmic trading oversight. First, the demonstrated ability of transformer models to consistently outperform traditional strategies highlights the growing role of advanced machine learning in generating high risk-adjusted returns. This may prompt regulatory bodies to consider new guidelines for algorithmic trading practices, especially regarding transparency and risk management. Furthermore, given the long out-of-sample testing period and robust methodology employed, these findings may encourage policy discussions around implementing stricter standards for the validation and monitoring of algorithmic models to safeguard against overfitting and ensure consistent performance. Finally, as these advanced models could widen the gap between retail and institutional investors, policies may be required to promote equitable access to AI-driven trading technologies.